Maxine:

There are 24 pictures on a roll of film. Adelle had 84 pictures from her trip home. She said that 12 other pictures were spoiled. How many rolls of film did Adelle use? A. 3 rolls B. 4 rolls C. 6 rolls D. 8 rolls

kylenickel4:

she would have used 4 rolls

Maxine:

how?

kylenickel4:

Adelle took 84+12, or 96 pictures in all, including the unspoiled ones too. And since there are 24 pictures in each roll, you divide 96 by 24. 96/24 gives you 4. So Adelle used 4 rolls of film.

emilybethch:

oop im too late sorry

Maxine:

it ok thanks kylenickle4

bridgetmadle5:

Find the total pictures taken: Total pictures that turns out good = 84 Total pictures that turns out spoilt = 12 Total pictures taken = 84 + 12 = 100 Find the number of rolls used: 1 roll = 24 pictures Number of rolls = 100 ÷ 24 = 4.17 ⇒ Adelle has used a little more than 4 films ⇒ Adelle has used 5 rolls of films

kylenickel4:

Thats how I am really good at these types of problems